OpenStudy (anonymous):

Another shopkeeper bought 100 pans at $5 each. He sold 63 at $6 each and x at $4 each. He did not sell all the pans nor enough to make an overall profit. Form an inequality in x. Hence find the greatest possible number of pans that were sold. Answer

OpenStudy (anonymous):

use the 6*63 +4x<= 100*55

OpenStudy (anonymous):

x+63 < =100

OpenStudy (anonymous):

to make a profit he would need more than $500 from sales. He already has 63*6= 378 this means that the total money he has made is less than 500 and the total money he has made is equal to 378 + 4x < 500 as $4 times x is the total money made from the cheaper pans. subtracting 378 from both sides gives 4x < 122 and now you simply need to divide by 4 to get the correct answer :)

OpenStudy (anonymous):

Remember to round down as he could not half sold half a pan and if you round up your total value will exceed what the inequality allows